A local council raises the price of car parking from E3 per day to {5 per day and finds that usage of car parks contracts from 1,200 cars a day to 900 cars per day.
Calculate the price elasticity of demand for this price change and calculate whether total revenue from the car park rises or falls
1
Expert's answer
2020-10-15T02:56:43-0400

Percentage change (price) is given by

"(5-3)=2"

"\frac{2}{3}×100"%"=66.67"%


Percentage change in demand is given by,

"(900-1200)= -300"


"\frac{-300}{1200}×100%"%"=-25"%


Price elasticity of demand is


"\frac{-25}{66.67}=-0.3750"


Therefore demand is price inelastic

Total revenue

Revenue at £ 3 per day is


"\u00a33\u00d71200=\u00a33600"


Revenue at £5 per day is


"\u00a35 \u00d7900=\u00a34500"


Therefore; revenue rises when the price elasticity of demand is less than one and this makes business raise their selling price.